OpenStudy (anonymous):

Is this right? please help me!! thank you <333 Simplify: 40v^2 over 35v^4 + 20v^3 over 5v my answer: 7v^4 over 2

OpenStudy (anonymous):

\[\frac{\frac{40v^2}{35v^4 + 20v^3}}{ 5v}\]\

OpenStudy (anonymous):

or is it an addition?

jimthompson5910 (jim_thompson5910):

could it be that the problem is actually \[\Large \frac{40v^2}{35v^4} + \frac{20v^3}{5v}\]

OpenStudy (anonymous):

yeah that is what i was wondering too

OpenStudy (anonymous):

you are correct, @jim_thompson5910

jimthompson5910 (jim_thompson5910):

\[\Large \frac{40v^2}{35v^4} + \frac{20v^3}{5v}\] \[\Large \frac{40v^2}{35v^4} + \frac{7v^3*20v^3}{7v^3*5v}\] \[\Large \frac{40v^2}{35v^4} + \frac{140v^6}{35v^4}\] see where to go from here?

OpenStudy (anonymous):

add both numirators?

jimthompson5910 (jim_thompson5910):

correct then simplify as much as possible

OpenStudy (anonymous):

180

OpenStudy (anonymous):

7/32?

jimthompson5910 (jim_thompson5910):

no you can't add 40v^2 to 140v^6 since they aren't like terms so you just leave it as \[\Large \frac{40v^2+140v^6}{35v^4}\] but you can still simplify this

OpenStudy (anonymous):

32/7

jimthompson5910 (jim_thompson5910):

this is because \[\Large \frac{40v^2+140v^6}{35v^4}\] becomes \[\Large \frac{20v^2(2+7v^4)}{35v^4}\] and that simplifies even further

OpenStudy (anonymous):

my final answer is 2 over 7v^4

jimthompson5910 (jim_thompson5910):

\[\Large \frac{40v^2+140v^6}{35v^4}\] \[\Large \frac{20v^2(2+7v^4)}{35v^4}\] \[\Large \frac{4(2+7v^4)}{7v^2}\] \[\Large \frac{8+28v^4}{7v^2}\]

OpenStudy (anonymous):

thats not an option....

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

what are your other options?

OpenStudy (anonymous):